OpenStudy (anonymous):

Given the points (2, 3) and (6, 8), write the point-slope form of the equation. A. y - 3 = x - 2 B. y - 3 = 5/4 (x - 2) C. y - 8 = 4/5 (x - 6)

OpenStudy (anonymous):

\[x-x _{1}/x_{0}-x_{1}=y-y_{1}/y_{0}-y_{1}\] is the formula to obtain a line from two points

OpenStudy (anonymous):

Can u just give me the answer. plz

OpenStudy (anonymous):

xD ok x-2/6-2=y-3/8-3 => x-2/4=y-3/5 => 5x-10=4y-12 => 5/4(x-2)=y-3 so the answer is B

OpenStudy (anonymous):

:) thnks

OpenStudy (anonymous):

you're welcome ;)

OpenStudy (anonymous):

k i have more for u but not rite now so dnt go far. lol ;)

OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line which passes through the points (1, 6) and (2, 9). A. y = 2x + 3 B. y = 3x + 3 C. y = 6x + 9 D. y = 3x + 6

OpenStudy (anonymous):

another time use that formula: x-1/2-1=y-6/9-6 => x-1=y-6/3 => 3x-3=y-6 => y=3x +3 answer is B ;) simple

OpenStudy (anonymous):

Which of the following is the equation for the horizontal line that passes through (-1, -3)? A. y = -1 B. y = -3 C. x = -1 D. x = -3

OpenStudy (anonymous):

horizontal line mean that have the same form of the x equation, so y=-3, -3 because this line pass from the -3 coordinate

OpenStudy (anonymous):

Last one.... Write the slope-intercept form of the equation of the line that passes through (4, 2) with a slope of 3. A. y = 3x - 10 B. y = 3x + 4 C. y - 2 = 3 (x - 4) D. y = 4x + 3

OpenStudy (anonymous):

slope is the m of the y=mx+q equation, so the equation will have the y=4x+q form... so what is the q? simply, sostitute the x and y with the coordinates of the point, so you have 2=3*4+q => q=2-12 => q=-10 so the equation will be y=3x-10

OpenStudy (anonymous):

I hope you have completely unterstand this exercise, made a graph will help a lot...

OpenStudy (anonymous):

thnks for ur help.

OpenStudy (anonymous):

for the science honor! XD